• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gaskell

    Log entry time 16:03:05 on December 09, 2008

    Entry number 252740

    keyword=Day shift summary

    Day Shift Summary

    8:00 Continue production running on 3He (pi+ transverse) from last shift. 8:15 Beam gone for BS. 9:00 In Controlled Access - lots of work ongoing. 11:30 HRS ramped and ready. 12:23 Back to beam permit. 13:09 Begin data taking on reference cell (H2)

    At end of shift, we are almost done with the 3 hours of H2 ref. cell running.

    Run # events Comment 4337 540k lower current due to INJ probs.adaq 4338 N/A test 4339 N/A test 4340 N/A ped. run 4341-4343 N/A test runs 4343 N/A spot++ run 4344 N/A short run - high dead time 4345 1 M 4346 1 M 4347 in progress